# Addressing a fundamental limitation in deep vision models: lack of spatial attention

### v1

@@ 3 A potential solution
The initial change map is set to all ones [^3] A similar initialization can be done by adding a blank frame at the beginning of the sequence.. At the frame level, frames are subtracted from each other ($numpy.abs(I_{t}-I_{t-1})$). Convolution is implemented by looping over spatial locations. If there is enough change (determined by L1 or L2 norm greater than threshold $\tau$) inside a receptive field (RF), that RF is processed; otherwise, it is discarded [^4] In practice, a layer uses its previously computed output at time $t-1$ and only updates some elements within it.. This results in significant computational savings, as the filters are not applied to unchanged locations. The conv layer keeps track of changes in its output map and generates a binary map where a 1 indicates a change. This map is sent to the subsequent pooling and convolution layers, and each layer saves its output for future use.
Notice that our implementation here is even slower than using nn.conv2d on CPU. The main point here is that sequential implementation on CPUs can be modified to save energy and to increase speed. Therefore, further investigation is needed to determine how this concept can be adopted for parallel processing on multi-core CUPs and GPUs.

### v2

@@ 2 Related work
Handling sparse data in deep learning, especially in convolutional neural networks (CNNs), often involves specialized architectures designed specifically for sparse inputs. One such approach is sparse convolutions, where the operations are limited only to non-zero elements, thereby skipping the unnecessary computation associated with dense convolutions. Submanifold Sparse Convolutions (SSC) [6], for instance, are particularly effective in maintaining sparsity by ensuring that only active (non-zero) features in each layer remain sparse throughout the network. This approach has proven useful in tasks like 3D object detection, where input data is inherently sparse (e.g., LiDAR point clouds). Another method is coordinate-based convolutions, which directly work with sparse coordinate inputs and avoid dense data representation altogether. Additionally, models like Sparse R-CNN [10] focus on sparse proposals to efficiently handle object detection by leveraging sparse inputs and predictions without requiring exhaustive feature maps. These methods are key in fields where data is naturally sparse, such as 3D vision and graph-based tasks, allowing CNNs to process information more efficiently by focusing only on relevant data points while maintaining performance.
@@ 3 A potential solution
The initial change map is set to all ones [^3] A similar initialization can be done by adding a blank frame at the beginning of the sequence.. At the frame level, frames are subtracted from each other ($np.abs(I_{t}-I_{t-1})$). Convolution is implemented by looping over spatial locations. If there is enough change (determined by L1 or L2 norm greater than threshold $\tau$) inside a receptive field (RF), that RF is processed; otherwise, it is discarded [^4] In practice, a layer uses its previously computed output at time $t-1$ and only updates some elements within it.. This results in significant computational savings, as the filters are not applied to unchanged locations. The conv layer keeps track of changes in its output map and generates a binary map where a 1 indicates a change. This map is sent to the subsequent pooling and convolution layers, and each layer saves its output for future use.
Three types of CNNs (Figure 3) are compared. CNN1 is the classic version that uses ‘nn.Conv2d‘. CNN2 sequentially applies convolution to image regions. CNN3 is similar to CNN2 but skips regions that have not changed. Notice that our implementation here (CNN3) is even slower than using nn.conv2d on CPU (CNN1). The main point here is that sequential implementation on CPUs can be modified to save energy and to increase speed. Therefore, further investigation is needed to determine how this concept can be adopted for parallel processing on multi-core CPUs and GPUs.
moving camera!!
@@ 4.2 Experiment II: Processing shifted versions of the same image
also mention if we have a tensor of size W x H x C channels (say C=100), and if your kernel size is KxK, then you save KxkxC operations are saved.
we were mainly concerned with inference time but maybe this can be also applied for training. finally applying this idea for other aritechtures and tasks such as face detecrtion (laptop camera etc) can be interesting.
@@ 5 Conclusion
Future work could explore techniques for handling sparse data, such as those discussed in the related works section. While this study focused on CPU optimization, extending these methods for GPU implementation is a promising direction for further research.